OpenStudy (anonymous):

What is the difference? Express your answer in lowest terms. - 2/11 - (- 6/11)

OpenStudy (isaiah.feynman):

\[a-(-b) = a+b\]

OpenStudy (anonymous):

???

OpenStudy (isaiah.feynman):

Also, if you have two fractions with the same denominator, sub\[\frac{ a }{ c } - \frac{ b }{ c } = \frac{ a - b }{ c }\]tracting goes

OpenStudy (anonymous):

A & B & C's ? What?

OpenStudy (isaiah.feynman):

I'm just giving a generic explanation for subtracting fractions.